### v2

@@ Appendix B Endogenous peer effects
for some (deterministic) $\varepsilon_{i}$ independent of $\mathbf{d}$. By (15), we can rewrite it as (recall that $\mathbf{L}$ defines the left-normalized Laplacian in Equation (15))
$Y_{i}(\mathbf{d})=\alpha+\beta_{n}\mathbf{d}_{i}+\kappa_{n}\sum_{j\neq i}\mathbf{L}_{i,j}\mathbf{d}_{j}+\gamma_{n}\sum_{j\neq i}\mathbf{L}_{i,j}Y_{i}(\mathbf{d})+\varepsilon_{i}.$
@@
$+\gamma_{n}\underbrace{\left\{\underbrace{\sum_{g=0}^{G-1}\gamma_{n}^{g}\mathbf{L}^{g+1}(\beta_{n}\mathbf{d}+\kappa_{n}\bar{\mathbf{d}}^{A})}_{\text{Effect of }G\text{ closest friends}}+\underbrace{\gamma_{n}^{G}\sum_{g=0}^{\infty}\gamma_{n}^{g}\mathbf{L}^{g+G+1}(\beta_{n}\mathbf{d}+\kappa_{n}\bar{\mathbf{d}}^{A})}_{\text{Approximation error from $G+1$ neighbors' effects}}\right\}}_{\text{Interference from higher order friends}}+\tilde{\epsilon}_{i},$
where $\alpha_{n}=(\alpha_{n,1},\ldots,\alpha_{n,n})^{\prime}=\alpha(\mathbf{I}-\gamma_{n}\mathbf{L})^{-1}\mathbf{1}$ and $\tilde{\epsilon}=(\tilde{\epsilon}_{1},\ldots,\tilde{\epsilon}_{n})^{\prime}=(\mathbf{I}-\gamma_{n}\mathbf{L})^{-1}\epsilon$.
Observe that the additional term that depends on higher order interference, multiplies by a factor $\gamma_{n}\beta_{n}+\gamma_{n}\kappa_{n}$. By letting each of these coefficients $\beta_{n},\kappa_{n},\gamma_{n}\propto\bar{\phi}_{n}$, Assumption 1 holds up to an order $o(1/n)$ (see Remark 2) for an adjacency matrix $\mathbf{A}_{G}$ where two individuals are connected if they are friends up to order $G$, provided that $\bar{\phi}_{n}^{G+1}\sum_{g=0}^{\infty}\bar{\phi}_{n}^{g}\mathbf{L}^{g+G+1}\mathbf{1}=o(1/n)$. This example illustrates that the approximation error due to local interference decreases exponentially fast in $G$ as $\bar{\phi}_{n}=o(1)$. However, in settings where researcher use an adjacency matrix $\mathbf{A}_{G}$ ($G>1$), we require stronger sparsity restrictions for our local asymptotic framework to hold (Assumption 3), with $\bar{\phi}_{n}G\mathcal{N}_{n,\max}^{2G}=o(1)$, where $\mathcal{N}_{n,\max}$ is the maximum degree under the original adjacency matrix $\mathbf{A}$.
@@
Assume that $\mathbf{A}$ is connected and $\epsilon_{1},\ldots,\epsilon_{n}$ are independent with $\mathrm{Var}(\epsilon_{i})=\sigma^{2}<\infty$. Then Assumption 7 holds if
@@
See Appendix C.15. ∎
Appendix C Proofs
@@ C.1 Auxiliary lemmas
Lemma C.1 (Zero covariances).
@@
Proof of Lemma C.1.
See Appendix C.3. ∎
Lemma C.1 states that two realized outcomes have zero covariance if two individuals (i) are in two different clusters, such that none of the two clusters contains a friend of the other individual, and (ii) are not friends or share a common friend (set), and if there is no friend of $j$ in a cluster that contains a friend of $i$. Lemma C.1 is equivalent to state that $\mu_{i}(D_{i},\mathbf{D}_{-i})[2D_{i}-1],\mu_{j}(D_{j},\mathbf{D}_{-j})[2D_{j}-1]$ have zero covariance if $B_{i}\cap B_{j}=\emptyset$. [^17] By definition, $B_{i}\cap B_{j}\neq\emptyset$ implies the existence of $i^{\prime}\in\{i\}\cup\mathcal{N}_{i},j^{\prime}\in\{j\}\cup\mathcal{N}_{j}$, and $k$ such that $c(i^{\prime})=c(k)=c(j^{\prime})$. Thus, $j^{\prime}\in B_{i}$. If $j^{\prime}=j$, then $j\in B_{i}$; otherwise, $j^{\prime}\in G_{i}$. For the converse, first note that $j\in B_{i}$ implies that $j\in B_{i}\cap B_{j}\neq\emptyset$ since $j\in B_{j}$ for any $j$. If $j\in G_{i}\setminus B_{i}$, then there exists $j^{\prime}\in\mathcal{N}_{j}$ such that $j^{\prime}\in B_{i}$. Since $j^{\prime}\in B_{j}$, $j^{\prime}\in B_{i}\cap B_{j}\neq\emptyset$. Next, we analyze the covariances for the remaining units.
Lemma C.2 (Non-zero Covariances).
@@
$\Big{|}\mathrm{Cov}\Big{(}\mu_{i}(D_{i},\mathbf{D}_{-i})\Big{[}2D_{i}-1\Big{]},\mu_{j}(D_{j},\mathbf{D}_{-j})\Big{[}2D_{j}-1\Big{]}\Big{)}\Big{|}=\mathcal{O}\Big{(}\bar{\phi}_{n}b_{n}(\mathcal{C}_{n})\Big{)}\quad\forall j:c(j)\neq c(i).$ (25)
@@
$=$ $\frac{1}{4}\Big{(}\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0})\Big{)}\Big{(}\mu_{j}(\mathbf{1})+\mu_{j}(\mathbf{0})\Big{)}+\mathcal{O}\Big{(}\bar{\phi}_{n}b_{n}(\mathcal{C}_{n})\Big{)}.$
@@
See Appendix C.4. ∎
Lemma C.2 characterizes the covariance between individuals in different clusters (Equation 25) and individuals in the same cluster (Equation 26). For individuals in different clusters, the covariance is of the same order of the bias, whereas for individuals in the same cluster, the covariance is $\mathcal{O}(1)$. The component $\frac{1}{4}\Big{(}\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0})\Big{)}\Big{(}\mu_{j}(\mathbf{1})+\mu_{j}(\mathbf{0})\Big{)}$ captures the covariance between individuals in the same clusters up-to a factor of order $\bar{\phi}_{n}b_{n}(\mathcal{C}_{n})$. The covariance between individuals in different clusters is zero if there are no individuals with neighbors in a different cluster since, in this case, the within-cluster covariance captures all the covariances between individuals.
C.2 Proof of Lemma 3.1
@@
$\sup_{\mu(1,\cdot)\in\otimes_{i=1}^{n}\mathcal{M}_{1,i}}\Big{|}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\Big{[}\mu_{i}(D_{i},\mathbf{D}_{-i})-\mu_{i}(\mathbf{1})|D_{i}=1\Big{]}\Big{|}$ $=\frac{\bar{\phi}_{n}}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}_{i}|}\mathbb{E}\Big{[}\sum_{k\in\mathcal{N}_{i}}(1-D_{k})\Big{|}|D_{i}=1\Big{]}$ (27)
@@
C.3 Proof of Lemma C.1
Let $C_{i}=\{c(v):v\in B_{i}\}$, and $G_{i}=\Big{\{}g\in\{1,\cdots,n\}:\mathcal{N}_{g}\cap B_{i}\neq\emptyset\Big{\}}$. Observe that we can write Lemma C.1 equivalently as stating that
@@
$\mathrm{Cov}\Big{(}\mu_{i}(1,\mathbf{D}_{-i})D_{i}-\mu_{i}(1,\mathbf{1})D_{i},\mu_{j}(1,\mathbf{D}_{-j})D_{j}\Big{)}$ $=\mathcal{O}\left(\mathbb{E}\Big{[}|\mu_{i}(1,\mathbf{D}_{-i})-\mu_{i}(1,\mathbf{1})|\Big{|}D_{i}=1\Big{]}\right).$
By (27) in the proof of Lemma 3.1, we have
$\mathbb{E}\Big{[}|\mu_{i}(1,\mathbf{D}_{-i})-\mu_{i}(1,\mathbf{1})|\Big{|}D_{i}=1\Big{]}=\frac{\bar{\phi}_{n}b_{n}(\mathcal{C}_{n})}{2}.$
Following a similar argument for the second component in Equation (30),
$\mathrm{Cov}\Big{(}\mu_{i}(1,\mathbf{1})D_{i},\mu_{j}(1,\mathbf{D}_{-j})D_{j}-\mu_{j}(1,\mathbf{1})D_{j}\Big{)}$ $=\mathcal{O}\left(\frac{\bar{\phi}_{n}}{2n}\sum_{j=1}^{n}\frac{1}{|\mathcal{N}_{j}|}\Big{|}i\in\mathcal{N}_{j}:c(j)\neq c(i)\Big{|}\right)$
$=\mathcal{O}\left(\bar{\phi}_{n}b_{n}(\mathcal{C}_{n})\right),$
@@
$\mathrm{Cov}\Big{(}\mu_{i}(1,\mathbf{D}_{-i})D_{i},\mu_{j}(1,\mathbf{D}_{-j})D_{j}\Big{)}=\mathcal{O}\left(\bar{\phi}_{n}b_{n}(\mathcal{C}_{n})\right).$
@@
$\mathrm{Cov}\Big{(}\mu_{i}(\mathbf{D})\Big{[}2D_{i}-1\Big{]},\mu_{j}(\mathbf{D})\Big{[}2D_{j}-1\Big{]}\Big{)}=\mathcal{O}\left(\bar{\phi}_{n}b_{n}(\mathcal{C}_{n})\right),$
@@
As a first step, we characterize the size of the set of units for which the covariances are not zero, but of order at most $\mathcal{O}(\bar{\phi}_{n}b_{n}(\mathcal{C}_{n}))$. Thus,
$\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}$ $=\frac{4}{n^{2}}\sum_{i,j}\mathrm{Cov}\Big{(}\mu_{i}(D_{i},\mathbf{D}_{-i})[2D_{i}-1],\mu_{j}(D_{j},\mathbf{D}_{-j})[2D_{j}-1]\Big{)}$
@@
$=\frac{4}{n^{2}}\sum_{i,j:c(i)=c(j)}\frac{1}{4}(\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0}))(\mu_{j}(\mathbf{1})+\mu_{j}(\mathbf{0}))+\mathcal{O}\left(\frac{\bar{\phi}_{n}b_{n}(\mathcal{C}_{n})}{n^{2}}\cdot\sum_{i,j:c(i)=c(j)}1\right)+$
@@
$\qquad\mathcal{O}\left(\frac{\bar{\phi}_{n}b_{n}(\mathcal{C}_{n})}{n^{2}}\cdot\sum_{i,j}I(j\in B_{i}\cup G_{i})\right)\qquad(\text{because }c(i)=c(j)\Longrightarrow j\in B_{i}\cup G_{i})$
$=\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}(\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0}))\right)^{2}+\mathcal{O}\left(\frac{\bar{\phi}_{n}b_{n}(\mathcal{C}_{n})}{n^{2}}\cdot\sum_{i}|B_{i}\cup G_{i}|\right).$
@@
where $C_{i}=\{c(v):v\in\{i\}\cup\mathcal{N}_{i}\}$ is defined in the proof of Lemma C.1. Thus,
@@
$\quad+\mathcal{O}(\bar{\phi}_{n}b_{n}(\mathcal{C}_{n}))\cdot\min\left\{\frac{1}{nK_{n}}\sum_{i}(|\mathcal{N}_{i}|+1)^{2},1\right\}.$ (33)
Since $\mathcal{N}_{n,\max}=\max\{\max_{i}|\mathcal{N}_{i}|,1\}$ and $b_{n}(\mathcal{C}_{n})\leq 1$,
@@
$\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}(\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0}))\right)^{2}+o(1/K_{n}).$ (36)
@@
$\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}\leq\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}\cdot\bar{\psi}+o(1/K_{n}).$
@@
Note that Assumption 3 implies $\bar{\phi}_{n}=o(1)$, we obtain that
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}\geq\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}\cdot\bar{\psi}+o(1/K_{n}).$

### v3

@@ B.5 Endogeneous peer effects
for some (deterministic) $\nu_{i}$. By (10), we can rewrite it as (recall that $\mathbf{L}$ defines the left-normalized Laplacian in Equation (10))
$Y_{i}(\mathbf{d})=\alpha+\beta_{n}\mathbf{d}_{i}+\kappa_{n}\sum_{j\neq i}\mathbf{L}_{i,j}\mathbf{d}_{j}+\gamma_{n}\sum_{j\neq i}\mathbf{L}_{i,j}Y_{i}(\mathbf{d})+\nu_{i}.$
@@
$+\gamma_{n}\underbrace{\left\{\underbrace{\sum_{g=0}^{G-1}\gamma_{n}^{g}\mathbf{L}^{g+1}(\beta_{n}\mathbf{d}+\kappa_{n}\bar{\mathbf{d}}^{A})}_{\text{Effect of }G\text{ closest friends}}+\underbrace{\gamma_{n}^{G}\sum_{g=0}^{\infty}\gamma_{n}^{g}\mathbf{L}^{g+G+1}(\beta_{n}\mathbf{d}+\kappa_{n}\bar{\mathbf{d}}^{A})}_{\text{Approximation error from $G+1$ neighbors' effects}}\right\}}_{\text{Interference from higher order friends}}+\tilde{\nu}_{i},$
where $\alpha_{n}=(\alpha_{n,1},\ldots,\alpha_{n,n})^{\prime}=\alpha(\mathbf{I}-\gamma_{n}\mathbf{L})^{-1}\mathbf{1}$ and $\tilde{\nu}=(\tilde{\nu}_{1},\ldots,\tilde{\nu}_{n})^{\prime}=(\mathbf{I}-\gamma_{n}\mathbf{L})^{-1}\nu$.
Observe that the additional term that depends on higher order interference, multiplies by a factor $\gamma_{n}\beta_{n}+\gamma_{n}\kappa_{n}$. Taking $\varepsilon=0$ in Assumption 3, so that spillovers are local to zero (but possibly non-negligible for inference), by letting each of these coefficients $\beta_{n},\kappa_{n},\gamma_{n}\propto\bar{\phi}_{n}$, Assumption 1 holds up to an order $o(1/n)$ (see Remark 2) for an adjacency matrix $\mathbf{A}_{G}$ where two individuals are connected if they are friends up to order $G$, provided that $\bar{\phi}_{n}^{G+1}\sum_{g=0}^{\infty}\bar{\phi}_{n}^{g}\mathbf{L}^{g+G+1}\mathbf{1}=o(1/n)$. This example illustrates that the approximation error due to local interference decreases exponentially fast in $G$ as $\bar{\phi}_{n}=o(1)$ is local to zero. However, in settings where researcher use an adjacency matrix $\mathbf{A}_{G}$ ($G>1$), we require stronger sparsity restrictions for our local asymptotic framework to hold (Assumption 3), with $\bar{\phi}_{n}G\mathcal{N}_{n,\max}^{2G}=o(1)$, where $\mathcal{N}_{n,\max}$ is the maximum degree under the original adjacency matrix $\mathbf{A}$.
@@
Assume that $\mathbf{A}$ is connected and $\nu_{1},\ldots,\nu_{n}$ are independent with $\mathrm{Var}(\nu_{i})=\sigma^{2}<\infty$. Then Assumption 7 holds if
@@
See Appendix E.8. ∎
B.6 Rates of convergence for Bernoulli and cluster design
Theorem B.7.
$\liminf_{n\rightarrow\infty}\frac{\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)}{\mathcal{B}_{n}^{*}(\mathcal{C}_{\mathrm{B},n},\lambda)}>1\quad\text{ if }\sqrt{K_{n}}\bar{\phi}_{n}\rightarrow 0\text{ and }K_{n}/n\leq\kappa,$ (24)
$\limsup_{n\rightarrow\infty}\frac{\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)}{\mathcal{B}_{n}^{*}(\mathcal{C}_{\mathrm{B},n},\lambda)}<1\quad\text{ if }\sqrt{K_{n}}\bar{\phi}_{n}\rightarrow\infty,\text{ and }b_{n}(\mathcal{C}_{n})\leq\delta,$
where $b_{n}(\mathcal{C}_{n})$ is defined in (6).
Proof.
See Appendix E.9. ∎
Theorem B.7 compares a Bernoulli design and a cluster design. Theorem B.7 states that we should not run a cluster experiment if the size of the spillover effects $\bar{\phi}_{n}$ goes to zero at an order faster than $1/\sqrt{K_{n}}$, where $K_{n}$ denotes the number of clusters, and the number of clusters is sufficiently smaller than the sample size (at an arbitrary rate). We should instead run a cluster design if spillover effects are larger in magnitude than $1/\sqrt{K_{n}}$. The assumption that $b_{n}(\mathcal{C}_{n})\leq\delta$ is equivalent to assuming that $\sum_{i=1}^{n}\frac{1}{|\mathcal{N}_{i}|}|\mathcal{N}_{i}\bigcap\{j:c(j)\neq c(i)\}|\leq\delta$, for some $\delta\in[0,1)$, i.e., the clustering decreases the bias of the Bernoulli design (since $b_{n}(\mathcal{C}_{\mathrm{B},n})=1$).
Table 4: Implications of Theorem B.7 and examples of rates.
Description Formulation Implication: Run Examples of Rates
small ##\## of clusters
very small ##\## of clusters
large ##\## of clusters
B.7 Estimated variance for inference on treatment effects
Returning to Section 6.1, in this section we discuss the estimation of the variance for ex-post inference. We define the ex post bias-aware confidence interval as
$\mathcal{I}_{\mathrm{ep}}=[\hat{\tau}\pm\chi(\bar{b}_{*,n},\hat{\sigma})],$ (25)
where $\hat{\sigma}(\mathcal{C}_{n})$ is an estimate of $\mathrm{Var}(\hat{\tau})$ using the experimental data.
Let
$X_{i}=2Y_{i}(2D_{i}-1),$ (26)
and
$B_{i}=\{v:\text{either }c(v)=c(i)\text{ or }c(v)=c(v^{\prime})\text{ for some }v^{\prime}\in\mathcal{N}_{i}\},\quad G_{i}=\{g:\mathcal{N}_{g}\cap B_{i}\neq\emptyset\}.$
Lemma D.1 below implies
$X_{j}\perp\!\!\!\perp X_{i},\quad\forall j\not\in B_{i}\cup G_{i}.$ (27)
It is easy to see that
$j\in B_{i}\cup G_{i}\Longleftrightarrow i\in B_{j}\cup G_{j}.$
Thus, (27) defines a undirected dependency graph among $(X_{1},\ldots,X_{n})$. Define
$i\sim j\text{ iff }j\in B_{i}\cup G_{i}.$
A particular choice is the network HAC estimator studied in Gao and Ding (2023), which is guaranteed to be positive.
$\hat{\sigma}^{2}(\mathcal{C}_{n})=\frac{1}{n^{2}}\sum_{i}\sum_{j\in B_{i}\cup G_{i}}K_{ij}(X_{i}-\hat{\tau})(X_{j}-\hat{\tau}),$ (28)
where $K=(K_{ij})_{i,j}$ is a PSD matrix satisfying
$K\succeq(I(i\sim j))_{i,j}$ (29)
For example, Gao and Ding (2023) suggests
$K=Q\max\{\Lambda,0\}Q^{T}$
where $Q\Lambda Q^{T}$ is the eigendecomposition of $(I(i\sim j))_{i,j}$. [^22] As with other work in the literature, Gao and Ding (2023) prove that $n\hat{\sigma}^{2}(\mathcal{C}_{n})\geq n\mathrm{Var}(\hat{\tau})+o_{\mathbb{P}}(1).$
Appendix C Additional numerical studies: Clustering with simulated networks
We report the estimated Pareto frontier of the proposed method and other clustering algorithms as a function of $\xi$ (weighting the variance and absolute bias). The proposed procedure uniformly leads to the lowest objective function. Note that because all other methods are Pareto dominated, then we should expect dominance also with respect to the sum of the variance and squared bias, here omitted for brevity only.
Appendix D Proofs
@@ D.1 Auxiliary lemmas
Lemma D.1 (Zero covariances).
@@
Proof of Lemma D.1.
See Appendix D.3. ∎
Lemma D.1 states that two realized outcomes have zero covariance if two individuals (i) are in two different clusters, such that none of the two clusters contains a friend of the other individual, and (ii) are not friends or share a common friend (set), and if there is no friend of $j$ in a cluster that contains a friend of $i$. Lemma D.1 is equivalent to state that $\mu_{i}(D_{i},\mathbf{D}_{-i})[2D_{i}-1],\mu_{j}(D_{j},\mathbf{D}_{-j})[2D_{j}-1]$ have zero covariance if $B_{i}\cap B_{j}=\emptyset$. [^23] By definition, $B_{i}\cap B_{j}\neq\emptyset$ implies the existence of $i^{\prime}\in\{i\}\cup\mathcal{N}_{i},j^{\prime}\in\{j\}\cup\mathcal{N}_{j}$, and $k$ such that $c(i^{\prime})=c(k)=c(j^{\prime})$. Thus, $j^{\prime}\in B_{i}$. If $j^{\prime}=j$, then $j\in B_{i}$; otherwise, $j^{\prime}\in G_{i}$. For the converse, first note that $j\in B_{i}$ implies that $j\in B_{i}\cap B_{j}\neq\emptyset$ since $j\in B_{j}$ for any $j$. If $j\in G_{i}\setminus B_{i}$, then there exists $j^{\prime}\in\mathcal{N}_{j}$ such that $j^{\prime}\in B_{i}$. Since $j^{\prime}\in B_{j}$, $j^{\prime}\in B_{i}\cap B_{j}\neq\emptyset$. Next, we analyze the covariances for the remaining units.
Lemma D.2 (Non-zero Covariances).
@@
$\Big{|}\mathrm{Cov}\Big{(}\mu_{i}(D_{i},\mathbf{D}_{-i})\Big{[}2D_{i}-1\Big{]},\mu_{j}(D_{j},\mathbf{D}_{-j})\Big{[}2D_{j}-1\Big{]}\Big{)}\Big{|}=\mathcal{O}(\bar{\phi}_{n})\quad\forall j:c(j)\neq c(i).$ (30)
@@
$=$ $\frac{1}{4}\Big{(}\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0})\Big{)}\Big{(}\mu_{j}(\mathbf{1})+\mu_{j}(\mathbf{0})\Big{)}+\mathcal{O}(\bar{\phi}_{n}).$
Above, the big-O terms are uniform over $i$ and $j$.
@@
See Appendix D.4. ∎
Lemma D.2 characterizes the covariance between individuals in different clusters (Equation 30) and individuals in the same cluster (Equation 31). For individuals in different clusters, the covariance is of the same order of the bias, whereas for individuals in the same cluster, the covariance is $\mathcal{O}(1)$. The component $\frac{1}{4}\Big{(}\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0})\Big{)}\Big{(}\mu_{j}(\mathbf{1})+\mu_{j}(\mathbf{0})\Big{)}$ captures the covariance between individuals in the same clusters up-to a factor of order $\bar{\phi}_{n}$. The covariance between individuals in different clusters is zero if there are no individuals with neighbors in a different cluster since, in this case, the within-cluster covariance captures all the covariances between individuals.
D.2 Proof of Lemma 3.1
@@
$\sup_{\mu(1,\cdot)\in\otimes_{i=1}^{n}\mathcal{M}_{1,i}}\Big{|}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\Big{[}\mu_{i}(D_{i},\mathbf{D}_{-i})-\mu_{i}(\mathbf{1})|D_{i}=1\Big{]}\Big{|}$ $\leq\frac{\bar{\phi}_{n}}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}_{i}|}\mathbb{E}\Big{[}\sum_{k\in\mathcal{N}_{i}}(1-D_{k})\Big{|}|D_{i}=1\Big{]}$ (32)
@@
D.3 Proof of Lemma D.1
Let $C_{i}=\{c(v):v\in B_{i}\}$, and $G_{i}=\Big{\{}g\in\{1,\cdots,n\}:\mathcal{N}_{g}\cap B_{i}\neq\emptyset\Big{\}}$. Observe that we can write Lemma D.1 equivalently as stating that
@@
$\mathrm{Cov}\Big{(}\mu_{i}(1,\mathbf{D}_{-i})D_{i}-\mu_{i}(1,\mathbf{1})D_{i},\mu_{j}(1,\mathbf{D}_{-j})D_{j}\Big{)}$ $=\mathcal{O}\left(|\mu_{i}(1,\mathbf{D}_{-i})-\mu_{i}(1,\mathbf{1})|\right)=\mathcal{O}(\bar{\phi}_{n}),$
where the last line is implied by Assumption 2. By symmetry, we also have
$\small\mathrm{Cov}\Big{(}\mu_{i}(1,\mathbf{1})D_{i},\mu_{j}(1,\mathbf{D}_{-j})D_{j}-\mu_{j}(1,\mathbf{1})D_{j}\Big{)}=\mathcal{O}\left(\bar{\phi}_{n}\right),$
@@
$\mathrm{Cov}\Big{(}\mu_{i}(1,\mathbf{D}_{-i})D_{i},\mu_{j}(1,\mathbf{D}_{-j})D_{j}\Big{)}=\mathcal{O}\left(\bar{\phi}_{n}\right).$
@@
$\mathrm{Cov}\Big{(}\mu_{i}(\mathbf{D})\Big{[}2D_{i}-1\Big{]},\mu_{j}(\mathbf{D})\Big{[}2D_{j}-1\Big{]}\Big{)}=\mathcal{O}\left(\bar{\phi}_{n}\right),$
@@
As a first step, we characterize the size of the set of units for which the covariances are not zero, but of order at most $\mathcal{O}(\bar{\phi}_{n})$. Thus,
$\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}$
$=\frac{4}{n^{2}}\sum_{i,j}\mathrm{Cov}\Big{(}\mu_{i}(D_{i},\mathbf{D}_{-i})[2D_{i}-1],\mu_{j}(D_{j},\mathbf{D}_{-j})[2D_{j}-1]\Big{)}$
@@
$=\frac{4}{n^{2}}\sum_{i,j:c(i)=c(j)}\frac{1}{4}(\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0}))(\mu_{j}(\mathbf{1})+\mu_{j}(\mathbf{0}))+\mathcal{O}\left(\frac{\bar{\phi}_{n}}{n^{2}}\cdot\sum_{i,j:c(i)=c(j)}1\right)+$
@@
$\qquad\mathcal{O}\left(\frac{\bar{\phi}_{n}}{n^{2}}\cdot\sum_{i,j}I(j\in B_{i}\cup G_{i})\right)\qquad(\text{because }c(i)=c(j)\Longrightarrow j\in B_{i}\cup G_{i})$
$=\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}(\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0}))\right)^{2}+\mathcal{O}\left(\frac{\bar{\phi}_{n}}{n^{2}}\cdot\sum_{i}|B_{i}\cup G_{i}|\right).$
@@
where $C_{i}=\{c(v):v\in\{i\}\cup\mathcal{N}_{i}\}$ is defined in the proof of Lemma D.1. Thus,
@@
$\quad+\mathcal{O}(\bar{\phi}_{n})\cdot\min\left\{\frac{1}{nK_{n}}\sum_{i}(|\mathcal{N}_{i}|+1)^{2},1\right\}.$ (38)
Since $\mathcal{N}_{n,\max}=\max\{\max_{i}|\mathcal{N}_{i}|,1\}$ ,
@@
$\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}(\mu_{i}(\mathbf{1})+\mu_{i}(\mathbf{0}))\right)^{2}+\mathcal{O}(\varepsilon/K_{n}).$ (41)
@@
$\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}\leq\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}\cdot\bar{\psi}+\mathcal{O}(\varepsilon/K_{n}).$
@@
Note that Assumption 3 implies $|\bar{\phi}_{n}|\leq\mathcal{O}(\varepsilon)$, we obtain that
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}\geq\frac{1}{n^{2}}\sum_{k=1}^{K_{n}}n_{k}^{2}\cdot\bar{\psi}+\mathcal{O}(\varepsilon/K_{n}).$